Everything is set for a second pumpdown tomorrow. We'll plan to start pumping after the 1pm meeting. Since the main volume is already at 12 torr, the roughing phase won't take nearly as long this time.
I've added new channels for the TP1 analog readings (current and speed) and for the two N2 tank pressure readings. Chub finished installing the new regulator and has run the transducer signal cable to the vacuum rack. In the morning he will terminate the cable and make the final connection to the Acromag.
Gautam and I updated the framebuilder config file, adding the newly-added channels to the list of those to be logged. We also set up a git repo containing all of the Python interlock/interfacing code: https://git.ligo.org/40m/vacpython. The idea is to use the issue tracker to systematically document any changes to the interlock code.